A baseball has 20 times the mass of a ping pong ball although the ping pong ball is moving.  twice as fast as the baseball. Compared to the kinetic energy of the ping pong ball, the baseball has.


1
Expert's answer
2021-06-14T08:18:31-0400

Let the kinetic energy of the ping pong ball be

"E_{ping-pong}=\frac{mv^2}{2}"


Then, the kinetic energy of the baseball equals

"E_{baseball}=\frac{20m\times(\frac{1}{2}v)^2}{2}=5\frac{mv^2}{2}" , or 5 times greater than the kinetic energy of a ping-pong ball.


Answer: "E_{baseball}=5E_{ping-pong}"